# Kernel Ridge Regression with Predicted Feature Inputs and Applications to Factor-Based Nonparametric Regression

### v1

@@ 3.4.3 Relating the local Rademacher complexity to the kernel complexity
By borrowing the results in Boucheron et al. (2000, 2003); Bartlett et al. (2005), we first establish the connection between the population and empirical local Rademacher complexities in Lemma 9 of Section B.3.1:
@@
provided that $\delta$ is not too small. Subsequently, we relate the empirical local Rademacher complexity $\widehat{\psi}_{x}(\delta)$ to the empirical kernel complexity $\widehat{R}_{x}(\delta)$ in (40). Both quantities depend on the predicted inputs $\widehat{g}(X_{1}),\ldots,\widehat{g}(X_{n})$. By following the argument used in the proof of Lemma 6.6 of Bartlett et al. (2005), we prove in Lemma 10 of Section B.3.2:
@@
with $\widehat{\mu}_{1}\geq\cdots\geq\widehat{\mu}_{n}$ being the eigenvalues of $\mathbf{K}$. Note that $\widehat{R}(\delta)$ can be regarded as the empirical counterpart of $R(\delta)$, and it is expected due to concentration that $\widehat{R}(\delta)$ and $R(\delta)$ have the same order. However, when the prediction error of $\widehat{g}(X_{i})$ is not negligible, it should manifest in the gap between $\widehat{R}_{x}(\delta)$ and $\widehat{R}(\delta)$. Our results in Lemmas 11 and 12 and Corollary 8 of Section B.3.3 characterize how the error of $\widehat{g}(X_{i})$ in predicting $Z_{i}$ inflates the empirical kernel complexity:
@@
The final step is to bound from above $\widehat{R}(\delta)$ by its population-level counterpart $R(\delta)$. This is proved in Lemma 13 of Section B.3.4:
@@
from which we finally derive the order of $\delta_{x}$ in Lemma 4 of Section B.2.2 as
@@
4 Application to factor-based nonparametric regression models
In this section, to provide a concrete application of our developed theory in Section 3, we consider the latent factor model (3) where the ambient features $X\in\mathcal{X}=\mathbb{R}^{p}$ are linearly related with the latent factor $Z\in\mathcal{Z}=\mathbb{R}^{r}$ with $r\ll p$. Under such model, it is reasonable to choose $\widehat{g}$ as a linear function to predict $Z$, as detailed in Section 4.1. In Section 4.2 we state the upper bound of the excess risk of the corresponding predictor while in Section 4.3 we prove a matching minimax lower bound, thereby establishing the minimax optimality of the proposed predictor.
4.1 Prediction of latent factors using PCA under model (3)
We first discuss the choice of $\widehat{g}$ to predict the latent factor $Z$. As mentioned in Remark 1, we construct such predictor from an auxiliary data $\mathbf{X}^{\prime}\in\mathbb{R}^{n^{\prime}\times p}$ (for simplicity, we assume $n^{\prime}=n$) that is independent of the training data $\mathcal{D}$. Under the factor model (3), the $n\times p$ matrix $\mathbf{X}^{\prime}=(X_{1}^{\prime},\ldots,X_{n}^{\prime})^{\top}$ satisfies
@@ 4.1 Prediction of latent factors using PCA under model (3)
$\begin{split}&(\widehat{\mathbf{Z}}^{\prime},\widehat{A})~{}=~{}\operatorname*{arg\,min}_{\mathbf{Z}^{\prime}\in\mathbb{R}^{n\times r},\ A\in\mathbb{R}^{p\times r}}~{}{1\over np}\bigl{\|}\mathbf{X}^{\prime}-\mathbf{Z}^{\prime}A^{\top}\bigr{\|}_{\text{F}}^{2},\\ &\textrm{subject to}\quad{1\over n}\mathbf{Z}^{{}^{\prime}\top}\mathbf{Z}^{\prime}=\mathbf{I}_{r},\quad{1\over p}A^{\top}A\textrm{ is diagonal}.\end{split}$
The above formulation uses $r$, the true number of latent factors. In the factor model literature, there exist several methods that are provably consistent for selecting $r$. See, for instance, Ahn and Horenstein (2013); Lam and Yao (2012); Bai and Ng (2002); Bing et al. (2022, 2021); Buja and Eyuboglu (1992). Write the singular value decomposition (SVD) of the normalized $\mathbf{X}^{\prime}$ as
${1\over\sqrt{np}}\mathbf{X}^{\prime}~{}=~{}\sum_{i=1}^{p}d_{i}u_{i}v_{i}^{\top}.$ (49)
Let $U_{r}=(u_{1},\ldots,u_{r})\in\mathbb{O}_{n\times r}$, $V_{r}=(v_{1},\ldots,v_{r})\in\mathbb{O}_{p\times r}$ and $D_{r}={\rm diag}(d_{1},\ldots,d_{r})$ containing the largest $r$ singular values. It is known (see, for instance, Bai (2003)) that the constrained optimization in (4.1) admits the following closed-form solution
$\widehat{A}=\sqrt{p}~{}V_{r}D_{r},\qquad\widehat{\mathbf{Z}}^{\prime}=\sqrt{n}~{}U_{r}=\mathbf{X}^{\prime}\left(V_{r}D_{r}^{-1}/\sqrt{p}\right)=:\mathbf{X}^{\prime}\widehat{B}^{\top}.$ (50)
@@
We state its excess risk bound in the next section, as an application of our theory in Section 3.
@@ 4.2 Upper bounds of the excess risk of $\widehat{f}\circ\widehat{B}$
From our theory in Section 3, prediction error of $\widehat{g}(X)=\widehat{B}X$ appears in the excess risk bound in the form of
$\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}=\bigl{\|}\Sigma_{Z}^{1/2}(\widehat{B}A-\mathbf{I}_{r})\bigr{\|}_{\text{F}}^{2}+\bigl{\|}\Sigma_{W}^{1/2}\widehat{B}^{\top}\bigr{\|}_{\text{F}}^{2}$ (51)
where we write $\Sigma_{Z}={\rm Cov}(Z)$ and $\Sigma_{W}={\rm Cov}(W)$. The second equality above is due to the independence between $Z$ and $W$. In view of (51), the matrix $\widehat{B}$ needs to ensure $\|\widehat{B}A-\mathbf{I}_{r}\|_{\text{F}}$ vanishing. Since this essentially requires to identify $A$, we adopt the following identifiability and regularity conditions that are commonly used in the factor model literature, see, for instance, Stock and Watson (2002); Bai (2003); Fan et al. (2013); Bai and Ng (2020). Throughout this section, we also treat the number of latent factors $r$ as fixed.
@@
The matrix $A$ and $\Sigma_{Z}$ satisfies $\Sigma_{Z}=\mathbf{I}_{r}$ and $p^{-1}A^{\top}A={\rm diag}(\lambda_{1},\ldots\lambda_{r})$ with $\lambda_{1},\ldots,\lambda_{r}$ being distinct. Moreover, there exists some absolute constant $0<c\leq C<\infty$ such that $\max_{j\in[p]}\|[\Sigma_{W}]_{j\cdot}\|_{1}<C$ and $c\leq\lambda_{r}\leq\lambda_{1}\leq C$.
@@
Both 6 and 7 are commonly assumed in the literature of factor models with the number of features allowed to diverge (Bai, 2003; Fan et al., 2013; Stock and Watson, 2002; Bai and Ng, 2006). The requirement of $\lambda_{1},\ldots,\lambda_{r}$ being distinct in (6) is used to ensure that the factor $Z$ can be consistently predicted. We remark that this requirement can be removed once kernels that are invariant to orthogonal transformation are used (see, also, Section 3.2.1).
Under Assumptions 6 and 7, the existing literature of approximate factor model (see, for instance, Bai and Ng (2020)) ensures that $\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}=\mathcal{O}_{\mathbb{P}}(n^{-1}+p^{-1}).$ Consequently, Theorem 2 together with $\mathbb{E}\Delta_{\widehat{g}}\lesssim\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}$ yields the following excess risk bounds of $\widehat{f}\circ\widehat{B}$.
@@
The term $1/p$ arises exclusively from using PCA to predict the latent factors under Assumptions 6 and 7. For linear kernels and $f_{*}\in\mathcal{H}_{K}$, we reduce to the factor-based linear regression models (2) and (3). Our results coincide with the excess risk bounds of PCR derived in Bing et al. (2021). When $f^{*}$ belongs to a hierarchical composition of functions in the Hölder class with some smoothness parameter $\gamma>0$, Fan and Gu (2024) derived the optimal prediction risk $(\log n/n)^{-{2\gamma\over 2\gamma+1}}+\log(p)/n+1/p$, achieved by regressing the response onto the leading PCs via deep neural nets. Corollary 5 on the other hand provides the excess risk bound of kernel ridge regression using the leading PCs under a broader function class. The obtained rate in the next section is further shown to be minimax optimal over a wide range of kernel functions.
@@ 4.3 Minimax lower bound of the excess risk
Regular kernels were already required in the past work of Yang et al. (2017) to derive the minimax lower bounds for estimating $f^{*}$ under the classical KRR setting with a fixed design. In our case, we further need the kernel function to be universal in order to characterize the effect of predicting $Z$. As mentioned in Remark 3, there is no approximation error for universal kernels. Our lower bounds below will hold pointwise for each $K\in\mathcal{K}$.
For the purpose of establishing minimax lower bounds, it suffices to choose $\epsilon\sim N(0,\sigma^{2})$ under model (4) and $(Z,W)$ being jointly Gaussian under model (3). For this setting, we use $\mathbb{P}_{\theta}$ to denote the set of all distributions of $\mathcal{D}$ as well as the test data point $(Y,X)$, parametrized by
@@
Let $\mathbb{E}_{\theta}$ denote its corresponding expectation. The following theorem states the minimax lower bounds of the excess risk under the above specification. Its proof can be found in Section B.9.
@@ 5 Extension to other loss functions
For a new data point $X\in\mathcal{X}$, we predict its corresponding response $Y$ by $(\widehat{f}\circ\widehat{g})(X)$. Our target of this section is to bound from above the excess risk of $\widehat{f}\circ\widehat{g}$,
@@
In addition to Assumptions 13 in the main paper, our analysis for general loss functions requires the following two assumptions. Recall that $\mathcal{F}_{b}$ is given in (37).
@@
Recall the kernel-related latent error $\mathbb{E}\Delta_{\widehat{g}}$ from (18) and $\delta_{n}$ as the fixed point of $R(\delta)$ in (22). The following theorem provides non-asymptotic upper bounds of the excess risk of $\widehat{f}\circ\widehat{g}$ with $\widehat{f}$ given in (53).
@@
Its proof can be found in Section B.10. ∎

### v2

@@ 3.4.3 Relating the local Rademacher complexity to the kernel complexity
By borrowing the results in Boucheron et al. (2000, 2003); Bartlett et al. (2005), we first establish the connection between the population and empirical local Rademacher complexities in Lemma 9 of Section A.3.1:
@@
provided that $\delta$ is not too small. Subsequently, we relate the empirical local Rademacher complexity $\widehat{\psi}_{x}(\delta)$ to the empirical kernel complexity $\widehat{R}_{x}(\delta)$ in (35). Both quantities depend on the predicted inputs $\widehat{g}(X_{1}),\ldots,\widehat{g}(X_{n})$. By following the argument used in the proof of Lemma 6.6 of Bartlett et al. (2005), we prove in Lemma 10 of Section A.3.2:
@@
with $\widehat{\mu}_{1}\geq\cdots\geq\widehat{\mu}_{n}$ being the eigenvalues of $\mathbf{K}$. Note that $\widehat{R}(\delta)$ can be regarded as the empirical counterpart of $R(\delta)$, and it is expected due to concentration that $\widehat{R}(\delta)$ and $R(\delta)$ have the same order. However, when the prediction error of $\widehat{g}(X_{i})$ is not negligible, it should manifest in the gap between $\widehat{R}_{x}(\delta)$ and $\widehat{R}(\delta)$. Our results in Lemmas 11 and 12 and Corollary 8 of Section A.3.3 characterize how the error of $\widehat{g}(X_{i})$ in predicting $Z_{i}$ inflates the empirical kernel complexity:
@@
The final step is to bound from above $\widehat{R}(\delta)$ by its population-level counterpart $R(\delta)$. This is proved in Lemma 13 of Section A.3.4:
@@
from which we finally derive the order of $\delta_{x}$ in Lemma 4 of Section A.2.2 as
@@
4 Application to factor-based regression models
To provide a concrete application of our developed theory in Section 3, we consider the factor model (3) where the observable features $X\in\mathcal{X}=\mathbb{R}^{p}$ are linearly related with the latent factor $Z\in\mathcal{Z}=\mathbb{R}^{r}$ with $r\ll p$. By centering, we can assume $Y$, $X$ and $Z$ in models (3) and (4) have zero mean. Under such model, it is reasonable to choose $\widehat{g}$ as a linear function to predict $Z$, as detailed in Section 4.1. In Section 4.2 we state the excess risk bound of the corresponding predictor while in Section 4.3 we prove a matching minimax lower bound, thereby establishing the minimax optimality of the proposed predictor.
4.1 Prediction of latent factors using PCA
We first discuss the choice of $\widehat{g}$ for predicting the latent factor $Z$. As mentioned in Remark 1, we construct such predictor from an auxiliary data $\mathbf{X}^{\prime}\in\mathbb{R}^{n^{\prime}\times p}$ (for simplicity, we assume $n^{\prime}=n$) that is independent of the training data $\mathcal{D}$. Under the factor model (3), the $n\times p$ matrix $\mathbf{X}^{\prime}=(X_{1}^{\prime},\ldots,X_{n}^{\prime})^{\top}$ satisfies
@@ 4.1 Prediction of latent factors using PCA
$\begin{split}&(\widehat{\mathbf{Z}}^{\prime},\widehat{A})~=~\operatorname*{arg\,min}_{\mathbf{Z}^{\prime}\in\mathbb{R}^{n\times r},\ A\in\mathbb{R}^{p\times r}}~{1\over np}\bigl\|\mathbf{X}^{\prime}-\mathbf{Z}^{\prime}A^{\top}\bigr\|_{\text{F}}^{2},\\ &\textrm{subject to}\quad\mathbf{Z}^{{}^{\prime}\top}\mathbf{Z}^{\prime}\textrm{ is diagonal},\quad A^{\top}A=p~\mathbf{I}_{r}.\end{split}$ (44)
The above formulation uses $r$, the true number of latent factors. In the factor model literature, there exist several methods that are provably consistent for selecting $r$. See, for instance, Ahn and Horenstein (2013); Bai and Ng (2002); Bing et al. (2022, 2021). Write the singular value decomposition (SVD) of the normalized $\mathbf{X}^{\prime}$ as
$\frac{1}{\sqrt{np}}\mathbf{X}^{\prime}~=~\sum_{i=1}^{p}d_{i}u_{i}v_{i}^{\top}.$
Let $U_{r}=(u_{1},\ldots,u_{r})\in\mathbb{O}_{n\times r}$, $V_{r}=(v_{1},\ldots,v_{r})\in\mathbb{O}_{p\times r}$ and $D_{r}={\rm diag}(d_{1},\ldots,d_{r})$ containing the largest $r$ singular values. It is known (see, for instance, Bai (2003)) that the constrained optimization in (44) admits the following closed-form solution
$\widehat{A}=\sqrt{p}~V_{r},\qquad\widehat{\mathbf{Z}}^{\prime}=\sqrt{n}~U_{r}D_{r}=\mathbf{X}^{\prime}(\widehat{A}/p)=:\mathbf{X}^{\prime}\widehat{B}^{\top}.$ (45)
@@
Its excess risk bound is stated in the next section, as an application of our theory in Section 3.
@@ 4.2 Upper bounds of the excess risk of $\widehat{f}\circ\widehat{B}$
From our theory in Section 3, by writing $\Sigma_{Z}={\rm Cov}(Z)$ and $\Sigma_{W}={\rm Cov}(W)$, prediction error of $\widehat{g}(X)=\widehat{B}X$ appears in the excess risk bound in the form of
$\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}=\|\Sigma_{Z}^{1/2}(\widehat{B}A-\mathbf{I}_{r})\|_{\text{F}}^{2}+\|\Sigma_{W}^{1/2}\widehat{B}^{\top}\|_{\text{F}}^{2}.$ (46)
The second equality is due to the independence between $Z$ and $W$. The matrix $\widehat{B}$ must therefore ensure that $\|\widehat{B}A-\mathbf{I}_{r}\|_{\text{F}}$ vanishes. Since this essentially requires identifying $A$, we adopt the following identifiability and regularity conditions commonly used in the factor model literature (see, for instance, Bai (2003); Fan et al. (2013); Bai and Ng (2020)). Throughout this section, we also treat the number of latent factors $r$ as fixed.
@@
The matrix $A$ and $\Sigma_{Z}$ satisfy $p^{-1}A^{\top}A=\mathbf{I}_{r}$ and $\Sigma_{Z}={\rm diag}(\lambda_{1},\ldots\lambda_{r})$ with $\lambda_{1},\ldots,\lambda_{r}$ being distinct. Moreover, there exist some absolute constants $0<c\leq C<\infty$ such that $\max_{j\in[p]}\|[\Sigma_{W}]_{j\cdot}\|_{1}<C$ and $c\leq\lambda_{r}\leq\lambda_{1}\leq C$.
@@
Both 6 and 7 are commonly assumed in the literature of factor models with the number of features allowed to diverge (Bai, 2003; Fan et al., 2013). The requirement of $\lambda_{1},\ldots,\lambda_{r}$ being distinct in (6) ensures that $Z$ can be consistently predicted. Such requirement can be dropped when using kernels invariant under orthogonal transformations (see also Section 3.2.1).
Under Assumptions 6 and 7, the existing literature (see, for instance, Bai and Ng (2020)) ensures that $\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}=\mathcal{O}_{\mathbb{P}}(n^{-1}+p^{-1}).$ Consequently, Theorem 1 together with $\mathbb{E}\Delta_{\widehat{g}}\lesssim\mathbb{E}\|Z-\widehat{B}X\|^{2}_{2}$ and Eq. 46 yields the following excess risk bounds of $\widehat{f}\circ\widehat{B}$.
@@
The term $1/p$ arises exclusively from using PCA to predict the latent factors under Assumptions 6 and 7. For linear kernels and $f_{*}\in\mathcal{H}_{K}$, we reduce to the factor-based linear regression models (2) and (3). Our results coincide with the excess risk bounds of PCR derived in Bing et al. (2021). When $f^{*}$ belongs to a hierarchical composition of functions in the Hölder class with some smoothness parameter $\gamma>0$, Fan and Gu (2024) derived the optimal prediction risk $(\log n/n)^{-{2\gamma\over 2\gamma+1}}+\log(p)/n+1/p$, achieved by regressing the response onto the leading PCs via deep neural nets. Corollary 5 on the other hand provides the excess risk bound of kernel ridge regression using the leading PCs under a broader function class. The obtained rate in the next section is shown to be minimax optimal over a wide range of kernel functions.
@@ 4.3 Minimax lower bound of the excess risk
Regular kernels were already required in the past work of Yang et al. (2017) to derive the minimax lower bounds for estimating $f^{*}$ under the classical KRR setting with a fixed design. In our case, we further need the kernel function to be universal in order to characterize the effect of predicting $Z$. As mentioned in Remark 3, there is no approximation error for universal kernels under mild assumptions. Our lower bounds below will hold pointwise for each $K\in\mathcal{K}$.
For the purpose of establishing minimax lower bounds, it suffices to choose $\epsilon\sim N(0,\sigma^{2})$ under model (4) and $(Z,W)$ being jointly Gaussian under model (3). For this setting, we use $\mathbb{P}_{\theta}$ to denote the set of all distributions of $\mathcal{D}$, parametrized by
@@
Let $\mathbb{E}_{\theta}$ denote its corresponding expectation. The following theorem states the minimax lower bounds of the excess risk under the above specification.
@@ 5 Extension to other loss functions
For a new data point $X\in\mathcal{X}$, we predict its corresponding response $Y$ via $(\widehat{f}\circ\widehat{g})(X)$. Our target of this section is to bound from above the excess risk of $\widehat{f}\circ\widehat{g}$,
@@
In addition to Assumptions 13, our analysis for general loss functions requires the following two assumptions. Recall that $\mathcal{F}_{b}$ is given in (32).
@@
Recall the kernel-related latent error from (18) and $\delta_{n}$ as the fixed point of $R(\delta)$ in (7). The following theorem provides non-asymptotic upper bounds of the excess risk of $\widehat{f}\circ\widehat{g}$ with $\widehat{f}$ given in (48).
@@
Its proof can be found in Section A.10. ∎